OpenStudy (anonymous):

Given the domain values { -2, 0, 2 }, find the range values for the function y = 3x - 2.

OpenStudy (helder_edwin):

replace x for the values -2, 0 and 2

OpenStudy (helder_edwin):

for example for x=2 then y=3(2)-2=6-2=4

OpenStudy (anonymous):

what??

OpenStudy (helder_edwin):

x=2 is mapped to y=4

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

try it yourself. with x=-2

OpenStudy (anonymous):

i dont understand....

OpenStudy (helder_edwin):

u have y=3x-2

OpenStudy (helder_edwin):

plug-in x=-2

OpenStudy (anonymous):

so y=-6-2?

OpenStudy (helder_edwin):

and -6-2=

OpenStudy (anonymous):

-8

OpenStudy (helder_edwin):

great this means that x=-2 maps into y=-8 what about x=0?

OpenStudy (anonymous):

-3

OpenStudy (helder_edwin):

no.

OpenStudy (helder_edwin):

y=3(0)-2=

OpenStudy (anonymous):

-2?

OpenStudy (helder_edwin):

yes.

OpenStudy (helder_edwin):

so the range is the set of y-values you got

OpenStudy (anonymous):

oh thanks :)

OpenStudy (helder_edwin):

u r welcome